OpenStudy (anonymous):

A truck is traveling at 45km/h for 2.0 hours due north, stops for 1.0 hour, and then heads due south at 75km/h for 3.0 hours. what is the average velocity of the car in km/h?

OpenStudy (shane_b):

\[V_{avg}=\frac{displacement}{time}\]\[V_{avg}=\frac{(45km/h*2h)-(75km/h*3h)}{6h}=?\]

OpenStudy (anonymous):

would the variable h be total hours?

OpenStudy (shane_b):

It's not a variable...it's just a unit h=hours

OpenStudy (shane_b):

Removing the units to make it cleared would look like this:\[V_{avg}=\frac{(45*2)-(75*3)}{6}=?km/h\]

OpenStudy (anonymous):

oh okay i got it! thank you so much!

OpenStudy (shane_b):

Note that the answer is negative only because I chose to use North as a positive direction

OpenStudy (anonymous):

yeah i got -22.5 and it said it was wrong so would it be positive instead?

OpenStudy (shane_b):

Yea...maybe it only wanted the magnitude |-22.5|=22.5

OpenStudy (anonymous):

still coming up wrong

OpenStudy (shane_b):

Not sure why...this is a very simple problem. The car ended up going 90km north and then 225km south. Therefore its total displacement from the starting point is 225-90=135km...and it did it over a 6 hour period. 135/6=22.5

OpenStudy (anonymous):

thanks